OpenStudy (anonymous):

A land developer sold 4 1/2 acres of his 15 1/5 acres. How many acres does he have left?

OpenStudy (ihelpyoulearn):

He has 10 7/1o left because 15 1/5-4 1/2=10 7/10! Glad to help anytime!

OpenStudy (bamboo888):

10 7/10 Because when you subtract, the denominators have to be the same. So to get the same, the only way you can go is up because 5 can't be divided by anything to get a whole number. So the denominators have to be 10. ***But what you do to the bottom, you must do to the top.*** so 15 1/5 = 15 2/10 and 4 1/2 = 4 5/10 then subtract. 15-4=11 BUT 2/10 is smaller than 5/10 so the 11 will become 10 and the fraction would be 7/10 so your final answer is 10 and 7/10.